OpenStudy (anonymous):

let g(x)= sqrt(8x)-sqrt(2x). find derivative of g(x)

OpenStudy (anonymous):

first write \[\sqrt{8x}=2\sqrt{2x}\] then get \[2\sqrt{2x}-\sqrt{2x}=\sqrt{2x}\] then take the derivative

OpenStudy (anonymous):

you get via the chain rule \[\frac{d}{dx}\sqrt{2x}=\frac{2}{2\sqrt{2x}}=\frac{1}{\sqrt{2x}}\]